The binding energy of solid copper metal is 448.7 kJ/mol. When Cu metal is illuminated with light of a wavelength exactly one-half of its photoelectric threshold wavelength, what is the kinetic energy of the ejected electrons, in eV? Write the answer to 3 significant digits. If no electrons are ejected, write -1 as the answer.
Answer #1

Wavelength of photon is half or threshold energy.

So, energy must be double.

E = 448.7*2 = 897.4 KJ/mol

KE of 1 mol of electron = E - Eo

= 897.4 KJ/mol - 448.7 KJ/mol

= 448.7 KJ/mol

KE of 1 electron = 448.7 / Avogadro’s number

= 448.7 / 6.022*10^23

= 7.451*10^-22 KJ

= 7.451*10^-19 J

= 7.451*10^-19 / 1.602*10^-19 eV

= 4.65 eV

Answer: 4.65 eV